OpenStudy (anonymous):

Correct 4 Medal ^.^ They have calculated all of the associated fees as well as any other expenses. They have $6,000 available to purchase the points, and plan to stay in the house for 14years. Loan A Loan B $278,100 $278,100 20-year fixed 20-year fixed 6.94% 6.62% 0 discount points 3 discount points M = $2,146.10 M=$2,093.13 ----------------- Loan A = 515064.0 Loan B = 780451.2 515064.0 -780451.2 = 26587.2

OpenStudy (anonymous):

They should purchase the discount points because they will be in the house long enough to justify the purchase. They should not purchase the discount points because they will not be in the house long enough to justify the purchase. They should not purchase the discount points because they do not have enough available cash. They should purchase the discount points because available cash should always be used to buy points.

OpenStudy (anonymous):

My answer : They should purchase the discount points because they will be in the house long enough to justify the purchase.

OpenStudy (maxdrizner):

loan A would be cheapest

OpenStudy (anonymous):

How I did this : ***12 * 20 =240** Loan A = 2146.10 * 240 + 0 = 515064.0 Loan B = 2093.13 * 240 + 278,100 = 780451.2

OpenStudy (maxdrizner):

what are you trying to find... how cheap it is or what the best deal is

OpenStudy (anonymous):

These are my choices : They should purchase the discount points because they will be in the house long enough to justify the purchase. They should not purchase the discount points because they will not be in the house long enough to justify the purchase. They should not purchase the discount points because they do not have enough available cash. They should purchase the discount points because available cash should always be used to buy points.

OpenStudy (anonymous):

My answer : They should purchase the discount points because they will be in the house long enough to justify the purchase.

OpenStudy (maxdrizner):

i think i found someone would would be able to help you but let me think cause i might be able to help

OpenStudy (maxdrizner):

that should be correct

OpenStudy (anonymous):

Thank you for the help :D !

OpenStudy (maxdrizner):

because if you do not have enough and you take out a loan, you will be in major dept and just everything will suck. never use credit cards either, only debit lol, these are my life lessons to you my friend

OpenStudy (maxdrizner):

and no problem

OpenStudy (maxdrizner):

and over time your property value will decrease unless you keep your house like spit shined lol

OpenStudy (anonymous):

What If I have 30 stray cats living with me :(

OpenStudy (maxdrizner):

well lol you will have to keep steaming the crud out of dem rugs lol. i rebuilt a playhouse and turned it into an office and there were piles of mouse crap i had to steam it

OpenStudy (anonymous):

I REBUILD A TREE HOUSE OUT OF TOOTH PICKS !!! until it stormed

OpenStudy (maxdrizner):

haha thats insane dude

OpenStudy (maxdrizner):

its pretty difficult to rebuild things when all you got is lumber laying around the yard and you are limited and all you have is a little hack saw and a hammer